A financial service recommends that a client purchase a corporate bond for $15,000 that earns 6% annual simple interest. how much additional money must be places in U.S. government securities that earn a simple interest rate of 3.5% so that the total annual interest earned from the two investments is 5% of the total investment?
Answer by Theo(13342)

15000 * .06 + x * .035 = (15000 + x) * .05
simplify this equation to get 15000*.06 + .035*x = .05*15000 + .05*x
simplify further to get 900 + .035*x = 750 + .05*x
subtract 750 from both sides of the equation and subtract .035*x from both sides of the equation to get 900 - 750 = .05*x - .035*x
simplify to get 150 = .015*x
divide both sides of the equation by .015 to get 10,000 = x

the client would need to purchase 10,000 additional dollars at 3.5% to get a combine investment of 5%.

15,000 * .06 + 10,000 * .035 = 900 + 350 = 1250.
1250 divided by a total investment of 15,000 + 10,000 = 25,000 is equal to 1250 / 25000 = .05 which is equal to 5%.